hello , can you please explain in detail, Thank you!
Answer #1

=100000+100000*1.05......
This is a Geometric Progression with first term (a) as 100000 and common ratio (r) as 1.05 and number of terms (n) as 35
We know sum of a GP of n terms with first term a and common ratio r=a*(1-r^n)/(1-r)
Hence, Total Salary earned
=100000*(1-1.05^35)/(1-1.05)
=9032030.735